find the LCM for 6, 28 AND 48
Found 2 solutions by stanbon, uma:
Answer by stanbon(75887)

find the LCM for 6,28,and 48
Express each one in prime factor form as follows:
6=2*3

28=2^2*7

48=2^4*3
-----------
Rule: The LCM is the product of each of the
prime factors present in its highest power.
So, you get:
2^4 * 3 *7 =336
That is the LCM
Cheers,
Stan H.

Answer by uma(370)

find the LCM for 6, 28 AND 48
solution:
6 = 2 * 3 ( one 2 and one 3)
28 = 2* 2 * 7 ( two 2's and one 7 )
48 = 2 * 2 * 2 * 2 * 3 * 1 ( four 2 's and one 3 )
LCM = 2*2*2*2*3*7
= 336